project evaluation. Revenue generated by a new fad product are as follows:

year 1 $40,000

year 2 $30,000

year 3 $20,000

year 4 $10,000

thereafter $0

Expenses are expected to be 40% of revenues and working capital required in each year is expected to be 20% of revenue in the following year. the product requires an immediate investment of $50,000 in plant and equipment.

a) if the opportunity cost of capital is 10%, what is the project's NPV?
Answer #1

Attaching the calculations and formulas used :-

с D E F G Particulars Year 1 Year 2 Year 3 Year 4 Year 5 Revenue 40,000 30,000 20,000 10,000 Expenses 16,000 12,000 8,000 4,0

B C Year 5 =0.4*G3 =0.2*H3 =G3-64-65 Particulars Year 1 Year 2 Year 3 Year 4 Revenue 40000 30000 20000 10000 Expenses =0.4*C3